bill533:

Sarah's age is rwice that of her sister. When you add Sarah's age to her sister's age, you get 33. How old is her sister? (a) Wirite an equation that represents the sistuation. Explain any variable used (b) Solve the equation from part (a). Show ur work. State your solution as a complete sentence

AngeI:

Ok so Sarah is x2 sisters age. So you would have to do 33-Sarahs age x2 to get the age of the sister

AngeI:

So honestly I dont know if this works all the time but for this I divided by 3

AngeI:

33/3=11 so like 11+11(2)=33 Meaning the sister is 11

bill533:

ok do i write that for a or b angel i get what you are saying

AngeI:

The equation for A

AngeI:

Would be a+a(2)=33

AngeI:

Then for B you would solve it

bill533:

oh wow thank you angel

AngeI:

No problem